OpenStudy (anonymous):

what is the slope and y-intercept of the equation 5x-3y=-6

OpenStudy (cgreenwade2000):

You have to set up your equation in slope intercept form which is: y=mx+b

OpenStudy (jackellyn):

put it in slope intercept form y=mx+b m is slope and b is y intercept.

OpenStudy (jackellyn):

Basically, you want to get y alone so you have to move everything to the other side. Example: 3x +2y= 8 2y = 8-3x ------> Subtract 3x from both sides y= 4-3/2x ------> Divide both sides by 2 y=-3/2x + 4 -----> slope intercept form m=-3/2 y=4 Try to do this with your equation.

OpenStudy (jackellyn):

What do you get @Bbjunkie ?

OpenStudy (anonymous):

5x-3y=-6 Solve the above for y.\[y=\frac{5 x}{3}+2 \]The slope is the coefficient of x and +2 is the y intercept.

OpenStudy (anonymous):

well the answers i have are a. slope=5/3; y-intercept=(0,2) b. slope=-5; y-intercept=(0,-6) c. slope=-3/5; y-intercept=(0,2) d. slope= 5/3; y-intercept= (0,-6)

OpenStudy (anonymous):

a. is the answer. Slope:\[\frac{5 x}{3}= \frac{5}{3}x \]The slope is 5/3 The y intercept is usually viewed as to where the line crosses the y axis. The x value of course is zero. (0,2) is a point with x=0 and y=2.

OpenStudy (anonymous):

Sorry about the response formatting. This site's LaTex implementation ain't no word processor.